OpenStudy (anonymous):

A box is constructed by cutting out square corners of a rectangler piece of cardboard and folding up the sides. If the cut out corners have sides with length x,then the volume of the box is given by the polynomial 4x^3-52x^2+160x. Find the volume of the box when x=4 inches.

OpenStudy (anonymous):

sub x=4 into 4x^3-52x^2+160x to get the volume.

OpenStudy (anonymous):

how would I do that

OpenStudy (anonymous):

4x^3-52x^2+160x = 4(4)^3-52(4)^2+160(4)

OpenStudy (anonymous):

is that the volume

OpenStudy (anonymous):

" the volume of the box is given by the polynomial 4x^3-52x^2+160x".....so....

OpenStudy (anonymous):

the volume is 160

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

no I am not sure what it is

OpenStudy (anonymous):

I think you are fully capable of doing scalar additions and multiplications....

OpenStudy (anonymous):

I can multiply and add but what am I adding and multiply I have never taken algebra before.

OpenStudy (anonymous):

you don't need algebra to calculate 4(4)^3-52(4)^2+160(4), there's no variables, did you notice? All constants, so just add and multiply like normal.

OpenStudy (anonymous):

16and 162

OpenStudy (anonymous):

4(4)^3-52(4)^2+160(4) = 4(64)-52(16)+160(4) = 256 - 832 + 640 = ?

OpenStudy (anonymous):

you should practice your addition and multiplication, you need to be proficient on basic matters such as those if you want to tackle more complex problem :D

OpenStudy (anonymous):

would it be 1206

OpenStudy (anonymous):

how did you get such a high number? May I ask what grade you are in?

OpenStudy (anonymous):

don"t be smart I am asking for help and you have to be nasty.I subtract 832 from286 and added 640.

OpenStudy (anonymous):

Nasty? In what way was I being nasty? I was concern that perhaps you are tackling questions normally beyond your grade level, or that your teacher is assigning questions that are much too difficult and thus ineffective in helping you learn the necessary materials that you should know before you go to the next grade. If you find offence in what I said, I hope it to be clearly offensive, but if it is something that could be otherwise, then you should ask me to clarify myself. In this case however, if you think you are uncomfortable in answering my question, which is somewhat involving your privacy to a certain degree then just simply so no, as I did use "May I ...".

OpenStudy (anonymous):

I just felt it to be offensive asking me what grade I am in. I have no idea how to do this problem but I can go elsewhere to figure it out.

OpenStudy (anonymous):

If you feel more comfortable getting help through a different medium then by all means please do. It is always better to learn in conditions that allow us to achieve the most in the time allowed. I find teachers to be really helpful, try asking them to help. Friends are great too, but I still prefer teachers, as they generally provide much more solid explanations and such. Good luck with your studies :D

OpenStudy (anonymous):

I am very sorry if I made you mad but I just got hurt that you ask that question

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem, and don't be shy to defend yourself if you feel you are violated in anyway.

OpenStudy (anonymous):

My teacher will not help me

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

Do you know of any sites that would help me

OpenStudy (anonymous):

Ah, sorry, I don't. I don't normally go on sites such as these, usually I ask teachers, friends and such :/